OpenStudy (anonymous):

A line has a slope of 9/4 and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct. A.-9x+4y=-56 B.9x+4y=56 C.y=9/4x-14 D.y=9/4x+14

OpenStudy (anonymous):

C. y=(9/4)x-14

OpenStudy (anonymous):

@robtobey there is more than just c all answer that are correct

OpenStudy (anonymous):

y = m x + b m = 9/4 and b = -14 y = 9/4 x + (-14) y = 9/4 x -14